OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? 4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7 I got y = 4x + 7

OpenStudy (helpblahblahblah):

your right

OpenStudy (anonymous):

its not one of the options

OpenStudy (helpblahblahblah):

ohh k let me try again becasue i tryied the problem and that what i got?

OpenStudy (anonymous):

at the most i think it could be D

OpenStudy (helpblahblahblah):

its D

OpenStudy (anonymous):

explain pls

OpenStudy (helpblahblahblah):

for sure because i put it in an x and y chart and i did the fomula on each pair and both answers were -7

OpenStudy (anonymous):

i didn't get the same asnwer, @Loser66

OpenStudy (anonymous):

y=4x+7 y-7=4x -7=4x-y

OpenStudy (anonymous):

Ohh so i can switch it around?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

but be careful with the signs

OpenStudy (anonymous):

it gets confusing

OpenStudy (anonymous):

4x - y = -7

OpenStudy (anonymous):

ty

OpenStudy (loser66):

y = 4x +7 minus y both sides 0 = 4x +7 -y -7 both sides -7 = 4x -y or 4x-y =-7 which is D